the average credit card debt for college seniors is 3262 if the debt is normally distributed with a standard deviation of 1100 find the probability that the senior owes between 2000 and 4000
Answer by stanbon(75887)

the average credit card debt for college seniors is 3262 if the debt is normally distributed with a standard deviation of 1100 find the probability that the senior owes between 2000 and 4000
-------------
z(2000) = (2000-3262)/1100 = -1.1473
z(4000) = (4000-3262)/1100 = 0.6709
----
P(2000< x <4000) = P(-1.1473< z <0.6709) = 0.6232 = 62.32%
===============
Cheers,
Stan H.